OpenStudy (anonymous):

Part 1. Use the 25° angle and the sine ratio to find the length of side b to two decimal places. I put: sin(25)=a/25 Is that right? Part 2. You can find the length of side a using cos 25° , tan 25° , or the Pythagorean theorem. Find the length of side a to 2 decimal places using one of these methods.

jimthompson5910 (jim_thompson5910):

Is there a picture of this?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

can you post it?

jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

so it should be sin(25) = b/25 and not sin(25) = a/25

jimthompson5910 (jim_thompson5910):

since in part 1 you're looking for b (and not a)

OpenStudy (anonymous):

what about part 2?

jimthompson5910 (jim_thompson5910):

part 2, you will have either... cos(25) = a/25 or tan(25) = b/a or a^2 + b^2 = 25 You'll use the value of b you found in part 1 to find the value of 'a' (keep in mind that a > 0) if you use method 2 or method 3.

jimthompson5910 (jim_thompson5910):

using cos(25) = a/25 is probably the easiest (in my opinion)

OpenStudy (anonymous):

Ok thank you so much man!

jimthompson5910 (jim_thompson5910):

you're welcome